For any irreversible process the net entropy change is Infinite Negative - Positive Zero A system ..... Is a region of consta please solve all
Answer #1

1) Answer = Positive

    Reason -

    According to the second law of thermodynamics " The total entropy of an isolated system that undergoes a change can never decrease". If the process is irreversible then the process will only occur with increase in entropy otherwise the process won't occur. Whereas in an reversible process, the total entropy of an isolated system remains constant.

Therefore, the change in entropy is always positive for an irreversible process. In general the total entropy and therefore the disorder always increase for an irreversible process.

2) Answer = Is a specified region, not neccessarily of constant volume, where transfer of energy and or mass is studied.

Reason -

A system is of 3 types-

i) Open system - Is a system where transfer of mass and energy both can occur.

ii) Closed system - Is a system where only transfer of energy across the boundaries of the system is allowed whereas transfer of mass is restricted.

iii) Isolated system - Is a system where both transfer of mass and energy is restricted.

Therefore a system is studied for the changes in mass and or energy taking place within the system.

3) Answer = Work

   Reason -

Work is not a thermodynamic property but it is a path function i.e. it depends upon the path followed that is initial state to final state. State functions are those functions which has specific value for each state i.e. they dont depend upon the path followed.

Whereas temperature, specific volume and pressure are the state functions. They have unique value for an given system. Hence they reflects the properties of the system.

4) Answer = Average Kinetic energy of the gaseous molecules.

Reason -

According to the average kinetic energy of gaseous molecules formula,

(1/2)mv2 = (3/2) kT

i.e. Average kinetic energy of gaseous molecules = (3/2) kT

Temperature is the function of kinetic energy of gaseous molecules. Higher the temperature higher is the kinetic energy of gaseous molecules.

5) Answer = Temperature is unaltered

    Reason-

   If we separate the term Isothermal i.e. Iso and thermal, meaning of Iso is same and that of thermal means temperature. Therefore meaning of isothermal process means the process taking place at constant temperature or same temperature.

So in Isothermal expansion of gases temperature is kept constant or is unaltered.